OpenStudy (anonymous):

Write a system of two equations using x and y for the following problem: The sum of two numbers is 170 and their difference is 26. Find the two numbers.

OpenStudy (anonymous):

x + y = 170 x - y = 26 Subtract them: 2y = 144, hence y = 72 then from second equation: x = 26 + y = 26 + 72 = 98